### v2

# JOINT OPTIMIZATION OF PIECEWISE LINEAR ENSEMBLES

Abstract
Tree ensembles achieve state-of-the-art performance on numerous prediction tasks. We propose Joint Optimization of Piecewise Linear Ensembles (JOPLEN), which jointly fits piecewise linear models at all leaf nodes of an existing tree ensemble. In addition to enhancing the expressiveness of an ensemble, JOPLEN allows several common penalties, including sparsity-promoting matrix norms and subspace-norms, to be applied to nonlinear prediction. We demonstrate the performance of JOPLEN on over 100 regression and classification datasets and with a variety of penalties. JOPLEN leads to improved prediction performance relative to not only standard random forest and gradient boosted tree ensembles, but also other methods for enhancing tree ensembles. We demonstrate that JOPLEN with a nuclear norm penalty learns subspace-aligned functions. Additionally, JOPLEN combined with a Dirty LASSO penalty is an effective feature selection method for nonlinear prediction in multitask learning.
Index Terms—  Joint, global, optimization, refinement, feature selection, subspace, ensemble, linear
1 Introduction
Ensemble methods combine multiple prediction functions into a single prediction function. A canonical example is a tree ensemble, where each tree forms a “partition” of the feature space, and each leaf is a “cell” of the partition with an associated constant value. Although neural networks have recently triumphed on structured data, tree ensembles, such as gradient boosting (GB), still achieve state-of-the-art performance on tabular (i.e., table-formatted) datasets across a broad range of applications [1].
Despite this success, standard implementations of tree ensembles are still limited, either by the piecewise constant fits at each cell, or by the greedy, suboptimal way in which the ensembles are trained. To address these limitations, several strategies to improve the performance of tree ensembles have been proposed. FASTEL [2] uses backpropagation to optimize the parameters of an ensemble of smooth, piecewise constant trees. Global refinement (GR) jointly refits all constant leaves of a tree ensemble after first running a standard training algorithm, such a random forests or gradient boosting [3]. Partition-wise Linear Models also perform joint optimization, but learn linear functions on axis-aligned and equally-spaced step functions [4]. Linear Forests (LF) [5] increase model expressiveness by replacing constant leaves with linear models.
Because tree ensembles are nonlinear and greedily constructed, it has also been challenging to incorporate structure-promoting penalties through joint optimization. RF feature selection typically relies on a heuristic related to the total impurity decrease associated to each feature. This approach may underselect correlated features, and requires further hueristics to extend to the multitask setting. GB feature selection [6] uses a greedy approximation of the $\ell_{1}$ norm, which penalizes the addition of new features to each subsequent tree. BoUTS [7] is a multitask extension that selects “universal” features (important for all tasks) and “task-specific” features by selecting features that maximize the minimum impurity decrease across all tasks. As these and other examples demonstrate, methods for incorporating feature sparsity and similar structural objectives into tree ensembles have been limited by the greedy nature of ensemble construction.
We propose Joint Optimization of Piecewise Linear Ensembles (JOPLEN), an extension of global refinement (GR) that applies GR to ensembles of piecewise linear functions. JOPLEN both ameliorates greedy optimization and increases model flexibility by jointly optimizing a hyperplane in each cell (e.g., leaf) in each partition (e.g., tree) of an ensemble. Besides improving performance, and unlike prior work on tree ensembles, JOPLEN is compatible with many standard structure-promoting regularizers, and thus provides a simple way to incorporate sparsity, subspace structure, and smoothness into nonlinear prediction. We demonstrate this capability on 149 real-world and synthetic datasets, and find that JOPLEN outperforms existing methods for regression, binary classification, and multitask feature selection, including GB, RF, LF, CatBoost, soft decision trees (FASTEL), and neural networks. Finally, we provide a GPU-accelerated Python package that is easily extensible to new loss functions and regularizers.
2 Methodology
To begin, we describe JOPLEN in the single task setting, and later extend it to multitask learning.
2.1 Single-task JOPLEN
Let $\mathcal{X}\doteq\mathbb{R}^{d}$ be the feature space $\mathcal{Y}$ the output space where, for regression, $\mathcal{Y}\doteq\mathbb{R}$, and for binary classification, $\mathcal{Y}\doteq\{-1,1\}$. Let $\{(\boldsymbol{x}_{n},y_{n})\}_{n=1}^{N}$ be a training dataset where $N\in\mathbb{N}$, $\boldsymbol{x}_{n}\in\mathcal{X}$, and $y_{n}\in\mathcal{Y}$. A model class $\mathcal{F}$ is a set of functions $f:\mathcal{X}\to\mathbb{R}$. Given $\mathcal{F}$, the goal of single-task supervised learning is to find an $f\in\mathcal{F}$ that accurately maps feature vectors to outputs. For regression and binary classification, predictions are made using $f(\boldsymbol{x})$ and $\operatorname*{\text{sign}}(f(\boldsymbol{x}))$, respectively.
JOPLEN is an instance of empirical risk minimization (ERM). In ERM, a prediction function $f$ is obtained as a solution of
$\hat{f}=\operatorname*{\arg\min}_{f\in\mathcal{F}}\sum_{n=1}^{N}\ell(y_{n},f(\boldsymbol{x}_{n}))+R(f)\;.$ (1)
where $R$ is a regularization term (penalty), and $\ell\colon\mathcal{Y}\times\mathbb{R}\to[0,\infty)$ is a loss function. We take $\ell$ to be the squared error loss for regression, and the logistic loss for binary classification.
We are interested in the setting where $f\in\mathcal{F}$ is an additive ensemble $f\doteq\sum_{p=1}^{P}f_{p}$, where $f_{p}$ is defined piecewise on a partition (e.g., a decision tree), and there are $P$ partitions. The partitions are fixed in advance, and may be obtained by running a standard implementation of a tree ensemble model. Further, we assume that $f_{p}(\boldsymbol{x})\doteq\sum_{c=1}^{C_{p}}(\langle\boldsymbol{w}_{p,c},\boldsymbol{x}\rangle+b_{p,c})\phi(\boldsymbol{x};p,c)$. Here ${C_{p}}$ is the number of “cells” in the partition (e.g., tree), $\boldsymbol{w}_{p,c}\in\mathbb{R}^{d}$ is a weight vector, $b_{p,c}\in\mathbb{R}$ is a bias term, and $\phi(\cdot;p,c)\colon\mathcal{X}\to\{0,1\}$ indicates whether data point $\boldsymbol{x}$ is in cell $c$ of partition $p$. Then, we denote a piecewise linear ensemble as
$f(\boldsymbol{x};\mathbf{W},\boldsymbol{b},\phi)=\sum_{p=1}^{P}\sum_{c=1}^{C_{p}}(\langle\boldsymbol{w}_{p,c},\boldsymbol{x}\rangle+b_{p,c})\phi(\boldsymbol{x};p,c)\;,$ (2)
where $C\doteq\sum_{p=1}^{P}{C_{p}}$ is the total number of cells, $\mathbf{W}\in\mathbb{R}^{C}$ is the matrix of all weight vectors, and $\boldsymbol{b}\in\mathbb{R}^{C}$ is a vector of all bias terms. JOPLEN is then defined by the solution of
$\operatorname*{\arg\min}_{\mathbf{W},\boldsymbol{b}}\frac{1}{N}\sum_{n=1}^{N}\ell\left(y_{n},f(\boldsymbol{x}_{n};\mathbf{W},\boldsymbol{b},\phi)\right)+R(\mathbf{W},\boldsymbol{b};\boldsymbol{x}_{n})\;,$ (3)
where $R$ is a regularization term (penalty). The notation $R(\mathbf{W},\boldsymbol{b};\boldsymbol{x}_{n})$ indicates that $R$ may depend on $\boldsymbol{x}_{n}$, but that only $\mathbf{W}$ and $\boldsymbol{b}$ may be penalized.
By running JOPLEN with a convex loss and regularization term, the overall objective is convex. We solve this problem using Nesterov’s accelerated gradient method, and use a proximal operator for any non-smooth penalties.
JOPLEN easily incorporates existing penalties, and provides a straightforward framework for extending well-known penalties for linear models to a nonlinear setting. Notable examples include sparsity-promoting matrix norms (e.g., $\ell_{p,q}$-norms [8, 9] and GrOWL [10]) and subspace norms (nuclear, Ky Fan [11], and schatten $p$-norms [12]). To illustrate this, in this section discuss two standard penalties: $\ell_{p,1}$-norms and the nuclear norm.
2.1.1 Single-task $\ell_{p,1}$-norm
Because we use linear leaf nodes, we can use sparsity-promoting penalties to preform feature selection. To learn a consistent sparsity pattern across all linear terms, we realize $R$ as an $\ell_{p,1}$ sparsity-promoting group norm (for $p\geq 1$),
$\begin{split}R_{p,1}(\mathbf{W},\boldsymbol{b};\boldsymbol{x}_{n})&\doteq\lambda\|\mathbf{W}\|_{p,1}\\ &\doteq\lambda\sum_{f=1}^{d}\left(\sum_{p,c=1}^{P,{C_{p}}}|w_{p,c}|^{p}\right)^{1/p}\;,\end{split}$ (4)
which leads to a row-sparse solution for $\mathbf{W}$. Concretely, we use the $\ell_{2,1}$ norm, which has a straightforward proximal operator [13, Eq. 6.8, p. 136].
2.1.2 Single-task nuclear norm
Given a singular value decomposition $\mathbf{W}\doteq\mathbf{U\Sigma V^{*}}$ for the conjugate transpose *, the nuclear norm is defined as
$R_{\text{nn}}(\mathbf{W},\boldsymbol{b};\boldsymbol{x}_{n})\doteq\lambda\|\mathbf{W}\|_{*}\doteq\lambda\sum_{i=1}^{\min(C,d)}|\mathbf{\Sigma}_{i,i}|\;.$ (5)
When applied to JOPLEN’s weights, the nuclear norm penalty empirically functions similarly to a group-norm penalty. While group-norms promote sparse solutions in an axis-aligned subspace, the nuclear norm apparently promotes “sparse” solutions along the “axes” of a data-defined subspace. This norm also has a straightforward proximal operator using singular value thresholding [13, Eq. 6.13].
2.2 Multitask JOPLEN
Suppose we have $T$ datasets indexed by $t\in\{1,...,T\}$: $\{(\boldsymbol{x}^{t}_{n},y^{t}_{n})\}^{N^{t}}_{n=1}$. Multitask JOPLEN is nearly identical to the single-task case, but includes an additional sum over tasks. Let $\mathbf{W}^{t}\in\mathbb{R}^{C^{t}\times d}$ for a task-specific number of cells $C^{t}$, $\boldsymbol{b}^{t}\in\mathbb{R}^{C^{t}},\phi^{t}\colon\mathcal{X}\to\{0,1\}$, and $\gamma^{t}\geq 0$ be the $t$-th task’s weight matrix, bias vector, partitioning function, and relative weight. For notational simplicity, define $\mathbf{W}\doteq[\mathbf{W}^{1}\cdots\mathbf{W}^{T}]$, $\mathbf{B}\doteq[\boldsymbol{b}^{1}\cdots\boldsymbol{b}^{T}]$, and $\boldsymbol{\gamma}\doteq[\gamma^{1}\cdots\gamma^{T}]^{\top}$. Then multitask JOPLEN is defined as
$\begin{split}\operatorname*{\arg\min}_{\mathbf{W},\mathbf{B}}\sum_{t=1}^{T}\frac{\gamma^{t}}{N^{t}}\sum_{n=1}^{N^{t}}&\ell\left(y_{n}^{t},f(\boldsymbol{x}_{n}^{t};\mathbf{W}^{t},\boldsymbol{b}^{t},\phi^{t})\right)\\ &+R(\mathbf{W},\mathbf{B};\boldsymbol{\gamma},\boldsymbol{x}_{n})\;,\end{split}$ (6)
and can be optimized using the same approach as in the single-task setting.
2.2.1 Dirty LASSO
JOPLEN with an extended Dirty LASSO (DL) [14] penalty performs feature selection for nonlinear multitask learning. Suppose that we know a priori that some features are important for all $T$ tasks, some features are important for only some tasks, and all other features are irrelevant. Then we can apply a JOPLEN extension of DL to perform “common” and “task-specific” feature selection.
DL decomposes the weight matrix $\mathbf{W}^{t}\doteq\mathbf{C}^{t}+\mathbf{T}^{t}$, encouraging $\mathbf{C}\doteq[\mathbf{C}^{1}\cdots\mathbf{C}^{T}]$ to be row-sparse (common features) and each $\mathbf{T}^{t}$ to be individually row-sparse (task-specific features), with potentially different sparsity patterns for each $\mathbf{T}^{t}$. Given penalty weights $\lambda_{\mathbf{C}},\lambda_{\mathbf{T}}\geq 0$, JOPLEN DL is
$\begin{split}\operatorname*{\arg\min}_{\mathbf{C},\mathbf{T},\mathbf{B}}\sum_{t=1}^{T}\Bigg{[}\frac{\gamma^{t}}{N^{t}}\sum_{n=1}^{N^{t}}&\ell\left(y_{n}^{t},f(\boldsymbol{x}_{n}^{t};\mathbf{C}^{t}+\mathbf{T}^{t},\boldsymbol{b}^{t},\phi^{t})\right)\\ &+\gamma^{t}\lambda_{\mathbf{T}}\|\mathbf{T}^{t}\|_{2,1}\Bigg{]}+\lambda_{\mathbf{C}}\|\mathbf{C}\|_{2,1}\;.\end{split}$ (7)
2.3 Laplacian regularization
The naïve use of piecewise-linear functions may lead to pathological discontinuities at cell boundaries. Thus, we utilize graph Laplacian regularization [15] to force nearby points to have similar values. In the single-task case,
$R_{\text{lap}}(\mathbf{W},\boldsymbol{b};\boldsymbol{x}_{n})\doteq\frac{\lambda}{2}\sum_{n_{1},n_{2}=1}^{N}\mathbf{K}(\boldsymbol{x}_{1},\boldsymbol{x}_{2})(\hat{y}_{1}-\hat{y}_{2})^{2}\;,$ (8)
where $\hat{y}_{i}\doteq f(\boldsymbol{x}_{i};\mathbf{W},\boldsymbol{b},\phi)$ is the model prediction for a feature vector $\boldsymbol{x}_{i}$ and $\mathbf{K}\colon\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ is a distance-based kernel. In this paper, $\mathbf{K}$ is a Gaussian radial basis function. Laplacian regularization can be naïvely applied to each task in the multitask setting as well.
3 Experiments
We evaluate JOPLEN in multiple regression and classification settings with several regularization terms.
3.1 Single-task regression and binary classification
We evaluate JOPLEN’s predictive performance on the “Penn Machine Learning Benchmark” (PMLB) [16]. The PMLB includes 284 regression, binary classification, and multiclass classification tasks. For simplicity, we focus on regression and binary classification. Some methods (e.g., GB) only handle real-valued features, so we drop all other features for such methods. We ignore datasets that have no real-valued features, or are too large for our GPU. This leaves 90 regression and 56 binary classification datasets. For each dataset, we perform a random 0.8/0.1/0.1 train/validation/test split.
To facilitate comparisons with previous methods [3], we jointly optimize partitions created using Gradient Boosted [17] and Random Forest [18] ensembles using CART trees.
We also evaluate JOPLEN with random Voronoi ensembles (Vor). Each partition in a Voronoi ensemble is created by sampling ${C_{p}}$ data points uniformly at random and creating Voronoi cells from these data points. This is to provide context for the performance of JOPLEN in Section 3.3 and Figure 3 b).
To demonstrate JOPLEN’s efficacy, we benchmark several alternative models: Gradient Boosted trees (GB), Random Forests (RF) [19], and Linear Forests (LF) [20, 5]; CatBoost [21]; Linear/Logistic Regression with a ridge penalty [19]; a neural network (NN) [22]; a differentiable tree ensemble (FASTEL) [2]; and global refinement (GR) [3]. GB, RF, and LF are baselines for demonstrating the performance improvement from joint optimization. CatBoost improves performance by correcting bias in GB’s functional gradient step [21]. FASTEL allows us to compare joint optimization and fully-differentiable ensembles. The provided implementation does not include classification losses, so we exclude it from the classification experiments. Ridge regression and NNs are standard approaches for modeling linear and non-linear functions, respectively. Finally, global refinement (GR) jointly optimizes all weights of a piecewise-constant tree ensemble. We exclude the global pruning step from [3] to more clearly demonstrate the direct performance improvement of JOPLEN’s piecewise-linear approach. These approaches provide a thorough comparison between JOPLEN and existing methods.
As in [3], we use the squared Frobenius norm of leaf nodes for both piecewise-constant and piecewise-linear ensembles. We also evaluate the Laplacian and the $\ell_{2,1}$-norm penalties (Sections 2.3 and 2.1.1).
We optimize model hyperparameters using Ax [23], an automated experiment platform that applies Bayesian and bandit optimization for continuous and discrete parameters, respectively. Ax uses 50 training/validation trials to model the structural risk as a function of hyperparameters, and we use the optimal hyperparameters to evaluate model performance on the test set. Hyperparameter ranges are documented in our code.
To facilitate plotting, we normalize the mean squared error (MSE) of each regression dataset by the MSE achieved by predicting the training mean and call this the “normalized” MSE. The 0/1 loss does not require normalization.
Fig. 1: Both plots show performance on the PMLB regression datasets. Each point is one data set, and is colored according to the original prediction method. “$\mathcal{L}$” and “F” indicate Laplacian + Frobenius (F) norm and F norm regularization. a) shows the normalized MSE on regression datasets. We do not plot datasets with an error above 1.5. The vertical dotted line indicates the naïve performance. $p$-values on the right compare the refitting method and with the original ensemble (e.g., “GB JOPLEN: $\mathcal{L}$” and “GB”). b) shows the 0/1 loss for classification datasets. For both plots, the black line indicates the median performance over all datasets.
Similar to [3], Figure 1 a) shows that both JOPLEN and GR significantly improve the performance of GB and RF on regression datasets using a one-sided Wilcoxon signed-rank test. JOPLEN also improves the performance of LF, and we find that JOPLEN with a GB partition outperforms all other regression models. Indeed, with Laplacian regularization, JOPLEN significantly outperforms CatBoost ($p=2.3\times 10^{-2}$) despite dropping all categorical features.
Although JOPLEN also improves performance for binary classification, the difference is less pronounced ($p=3.3\times 10^{-1}$ for GB). This decreased significance may be caused by the sample size (56 vs. 90 datasets), or the fact that the 0/1 loss only considers the sign of the prediction.
Interestingly, we observe that the partitioning method may significantly affect performance. In the regression setting, GB, LF, RF, and Voronoi model performances rank in this order. However, there is no significant difference in the classification setting. We suspect that there may be other partitioning methods that lead to further performance increases.
Finally, we note that JOPLEN outperforms the NN and FASTEL in general. Although the NNs and FASTEL perform very well on many datasets, they have a lower median performance because some datasets perform quite poorly. This is particularly true for FASTEL.
3.2 Single-task nuclear norm
In Figure 2 we demonstrate the effect of the nuclear norm on synthetic data. We draw 10,000 samples (100 are training data) uniformly on the square interval $[-1,1]^{2}$, and define $y_{n}\doteq f(\boldsymbol{x}_{n})$ where $f(\boldsymbol{x})\doteq\sin(\pi(x_{1}+x_{2}))+\varepsilon$ for an input feature $\boldsymbol{x}$ and $\varepsilon\sim\mathcal{N}(0,0.2)$. Here, we use random Voronoi ensembles and fix the number of partitions and cells to simplify visualization, and manually tune the penalty weights.
Fig. 2: a) and b) show the training and testing sets of a function that lies along a feature subspace. c) shows linear JOPLEN’s prediction using a Frobenius norm penalty, and d) shows the prediction using the nuclear norm penalty. The mean squared error (MSE) is reported below each method.
Notably, the nuclear norm causes the model to align its piecewise linear functions along the diagonal, forming a “consensus” across all linear models. By contrast, the Frobenius norm penalty allows each leaf to be optimized independently, and thus suffers from degraded performance.
3.3 Multitask Dirty LASSO
Next, we show that JOPLEN Dirty LASSO (DL) can achieve superior sparsity to the DL [14].
We selected two multitask regression datasets from the literature: “NanoChem” [7] and “SARCOS” [24]. “NanoChem” is a group of 7 small molecule, nanoparticle, and protein datasets that were collected to evaluate multitask feature selection performance [7]. These datasets have 1,205 physico-chemical features, and the sample sizes range from 127 to 11,079. The response variables are small molecule boiling point (1), Henry’s constant (2), logP (3), melting point (4), nanoparticle logP (5), zeta potential (6), and protein solubility (7). Similarly, “SARCOS” is a 7-task dataset that models the dynamics of a robotic arm [24], with 27 features and 48,933 samples for each task. These datasets were also split into train, validation, and test sets using a 0.8/0.1/0.1 ratio.
Using JOPLEN DL with tree ensembles will provide incorrect sparsity estimates; input features will be used to create each tree (not penalized) and then to train each linear leaf (penalized). In this case, simply analyzing the leaf weights will underestimate the number of features used by the JOPLEN model. To avoid this issue, we create our partitions using random Voronoi ensembles.
All penalties were manually tuned so that all methods achieve a similar test loss for each task. Because not all datasets are equally challenging to predict, we manually tune the $\gamma^{t}$ parameter for JOPLEN DL and the equivalent for BoUTS. No equivalent parameter exists in the community implementation of DL [25].
Fig. 3: a) demonstrates that JOPLEN DL learns common and task-specific features from the SARCOS dataset. The $x$-axis represents cells grouped by task, and the $y$-axis indicates the associated input feature. Blue indicates negative weights, red indicates positive weights. b) shows regression performance, with equal performance on the diagonal. JOPLEN’s performance is equal to or greater than that of DL (orange) on most tasks, and similar to that of BoUTS (teal).
Fig. 4: Comparison of the number of common and task-specific features selected by DL, BoUTS, and JOPLEN DL for the a) SARCOS and b) NanoChem datasets (fewer features is better).
We find that JOPLEN DL and BoUTS select significantly sparser feature sets than DL does, with JOPLEN providing the sparsest sets. In Figure 3 a) we see that joint optimization enables penalties to be shared across ensemble terms, providing structured sparsity. Further, Figures 4 and 3 b) show that JOPLEN selects far fewer features than DL approach while achieving similar or superior performance on all but one task (small molecule logP (3)). This is likely because of the significant disparity in the number of features selected for this task (102 for DL vs. 6 for JOPLEN).
4 Conclusions
The JOPLEN framework enables many different penalties to be incorporated into nonlinear methods. For example, it enables a straightforward approach to feature selection in nonlinear single- and multitask settings.
We find that JOPLEN beats global refinement (GR) and significantly outperforms other state-of-the-art methods on a wide variety of real-world and synthetic datasets. We demonstrate that Laplacian; Frobenius, nuclear, $\ell_{p,1}$-norm; and Dirty LASSO regulation have straightforward extensions using JOPLEN and improve model performance. Empirically, JOPLEN shines in settings where response variables is nonlinear but has structured sparsity, such as multitask feature selection. Indeed, JOPLEN achieves similar or higher feature sparsity and performance than BoUTS (Figure 3 b)) while using suboptimal Voronoi partitions (Figure 1). Future work may use BoUTS as a partitioner for JOPLEN to improve performance while maintaining a high degree of sparsity. Combining JOPLEN with global pruning or developing new partitioning methods may also lead to further improvements. Overall, these results suggest that JOPLEN is a promising approach to improving performance on tabular datasets.
We anticipate that JOPLEN will provide improvements in general regression and classification performance across a wide variety of fields. Additionally, we expect that JOPLEN’s piecewise linear formulation will lead to improved interpretability through increased feature sparsity. Finally, because JOPLEN allows the straightforward transfer of linear penalties to the nonlinear setting, we anticipate that it will greatly simplify the implementation of nonlinear penalized optimization problems (e.g., subspace-aligned nonlinear regression via the nuclear norm, Figure 2).
5 Code availability
A JAX (CPU and GPU compatible) version of our code is available on PyPI using pip install joplen. The source code is available at https://gitlab.eecs.umich.edu/mattrmd-public/joplen-repositories/joplen. The CuPY (GPU-only) implementation used in this paper is available via the tag mlsp_2024. The training, evaluation, and plotting code for this paper is available at https://gitlab.eecs.umich.edu/mattrmd-public/joplen-repositories/joplen-tests.